# Spatial Phonons: A Phenomenological Viscous Dark Energy Model for DESI

### v2

@@ 4 Cosmological Implications and Comparison with DESI
Using the benchmark value $H_{\star}/H_{0}=2.10$ we find,
$m_{\phi}\simeq\frac{2.10}{2\,\pi}\times H_{0}\simeq 0.334H_{0}$ (4.15)
The value of Hubble parameter today is $H_{0}\simeq 1.4\times 10^{-33}\,\text{eV}$ this gives,
$m_{\phi}\simeq 4.7\times 10^{-34}\,\text{eV}$ (4.16)
which is equivalent to,
$m_{\phi}=8.3\times 10^{-70}\ \mathrm{kg}$ (4.17)
The phonons are therefore extremely light. Their mass is only a few times the present Hubble scale, so their associated Compton wavelength is of order the cosmological horizon. This is consistent with their interpretation as infrared degrees of freedom that probe the large scale structure of the brane and respond to the global expansion rather than to local microscopic physics. In this picture the viscous response of the medium and the transient departure of $w_{\rm eff}(z)$ from minus one both originate from very soft collective excitations of space itself, which remain dynamically relevant only when the Hubble rate is comparable to the phonon mass gap and become negligible at much earlier and much later times.
@@ 5 Conclusion and Outlook
In this work we modeled cosmic acceleration as the effective dynamics of an elastic brane that represents space itself. The underlying geometric contribution is described by a Nambu–Goto type tension $T_{s}$ that survives as a residual cosmological constant in a minimal sequester framework. On top of this background tension we introduced a phonon fluid built from three scalar fields $\phi^{I}$ and an invariant $b=\sqrt{\det B_{IJ}}$ that controls the effective action $F(b)$. At the background level this construction reproduces a perfect fluid with energy density $\rho_{\rm ph}(b)$, pressure $p_{\rm ph}(b)$ and bulk modulus $K_{\rm ph}$. We parametrized the phonon sector by two dimensionless constants $\varepsilon$ and $\kappa$ which fix the ratio $(\rho_{\rm ph}+p_{\rm ph})/T_{s}$ and the bulk modulus in units of the space tension. This led to a simple and transparent relation for the phonon sound speed, $c_{s}^{2}=\kappa/\varepsilon$, with stability requiring $0<\kappa/\varepsilon\leq 1$.
@@
$w_{\rm eff}(H)=-1+\varepsilon-3\kappa\frac{x(H)}{1+x(H)^{2}},\qquad x(H)=H\tau(H)$
which makes it clear that a transient phantom phase can occur when $x(H)$ passes through unity, while $w_{\rm eff}$ tends to $1+\varepsilon$ at very early and very late times.
[…]
Several natural directions for future work emerge from this first exploration. The most important step is to derive the bulk viscosity $\zeta(H)$ and the relaxation time $\tau(H)$ from a more complete microscopic theory of the phonon sector. One would like to start from an explicit effective field theory for the longitudinal modes of the brane or for additional scalar degrees of freedom, compute their dispersion relations and interaction rates, and evaluate the stress tensor correlator that enters the Kubo formula. This program would allow one to determine $\zeta(H)$ and $\tau(H)$ from first principles rather than from phenomenology and would test whether the simple scaling used here is robust or only a useful approximation.
A second important step is to move beyond the homogeneous background and analyze cosmological perturbations in this framework. The phonon sound speed, the viscous pressure and any induced anisotropic stress could in principle affect the growth of density perturbations and the evolution of metric potentials. This makes it natural to confront the model with large scale structure observables such as redshift space distortion measurements and lensing data, in addition to background probes. At present it is not clear whether the perturbative effects of the viscoelastic sector would lead to observable deviations or remain hidden within current uncertainties. Only a detailed and consistent treatment of scalar perturbations can answer this question in a quantitative manner. Such an analysis would place bounds on $\kappa$, $\varepsilon$, the crossing scale $H_{\star}$ and the effective phonon mass, and would test whether the viscoelastic dark energy sector remains stable and well behaved at the level of fluctuations.
Finally, it would be interesting to embed the present phenomenological model into a broader picture of vacuum energy in gravity. The residual tension $T_{s}$ can be viewed as the leftover of a sequester mechanism that cancels matter induced vacuum energy and the phonon sector provides a concrete realization of dynamical dark energy on top of this geometric background. A more complete theory could clarify how the NG brane tension and the phonon fluid arise from a fundamental microscopic description and whether similar viscoelastic effects appear in other approaches to late time acceleration. The results presented here indicate that an elastic and viscous space is a viable candidate for the dark energy sector and they motivate a more systematic study of the microphysics and cosmological implications of such models.

### v3

@@ 4.10 Phonon Mass Scale associated with $H_{\star}$
Using the best fit value $h_{\star}=H_{\star}/H_{0}=1.794118$, this gives,
$m_{\phi}\;=\;\frac{h_{\star}}{2\pi}\,H_{0}\;\simeq\;0.28554\,H_{0}$ (4.22)
Adopting $H_{0}\simeq 1.4\times 10^{-33}\,\mathrm{eV}$, we obtain
$m_{\phi}\;\simeq\;4.00\times 10^{-34}\,\mathrm{eV}\;\simeq\;7.13\times 10^{-70}\,\mathrm{kg}$ (4.23)
This mass scale is around 30% of the present Hubble scale, so the associated Compton wavelength is of order the cosmological horizon. This is consistent with the interpretation of the phonons as infrared degrees of freedom that respond to the global expansion and become most relevant when $H$ is comparable to the mass gap scale.
@@ 5 Conclusion and Outlook
In this work we modeled cosmic acceleration as the effective dynamics of an elastic brane that represents space itself. The underlying geometric contribution is described by a Nambu–Goto type tension $T_{s}$ that survives as a residual cosmological constant in a minimal sequester framework. On top of this background tension we introduced a phonon fluid built from three scalar fields $\phi^{I}$ and an invariant $b=\sqrt{\det B_{IJ}}$ that controls the effective action $F(b)$. At the background level this construction reproduces a perfect fluid with energy density $\rho_{\rm ph}(b)$, pressure $p_{\rm ph}(b)$ and bulk modulus $K_{\rm ph}$. We parametrized the phonon sector by two dimensionless constants $\epsilon$ and $\kappa$ which fix the ratio $(\rho_{\rm ph}+p_{\rm ph})/T_{s}$ and the bulk modulus in units of the space tension. This led to a simple and transparent relation for the phonon sound speed, $c_{s}^{2}=\kappa/\epsilon$, with stability requiring $0<\kappa/\epsilon\leq 1$.
@@
$w_{\rm eff}(H)=-1+\epsilon-3\kappa\,\frac{x(H)}{1+x(H)^{2}}\,,\quad x(H)=H\,\tau(H)$ (5.1)
which makes it clear that a transient phantom phase can occur when $x(H)$ passes through unity, while $w_{\rm eff}$ tends to $1+\epsilon$ at very early and very late times.
Using a compressed Gaussian likelihood in the CPL plane derived from publicly released posterior chains for DESI DR1 BAO combined with Pantheon+ and Planck 2018, we find that the model can reproduce the central region of the published $(w_{0},w_{a})$ constraints without imposing any additional near luminal band beyond the basic causality prior $0<\kappa\leq\varepsilon$. In particular, the parameter scan naturally selects a best fit with $\kappa<\varepsilon<1$ and a small fractional splitting $(\varepsilon-\kappa)/\varepsilon\simeq 0.038$, which implies a near luminal sound speed $c_{s}^{2}=\kappa/\varepsilon\simeq 0.962$ as an output of the fit. This behaviour is physically natural in the present framework because the relevant phonon excitations are infrared collective oscillations of the brane with a mass gap $m_{\phi}=H_{\star}/(2\pi)$ set by a Hubble scale of order unity in redshift, so their Compton wavelength is of order the cosmological horizon. For such horizon scale modes, relativistic propagation is the generic expectation unless additional microphysical structure enforces a significantly smaller sound speed. In this sense, the compressed likelihood comparison does not merely accommodate near luminal phonons, it points to them in a particularly economical way.
Several natural directions for future work emerge from this first exploration. The most important step is to derive the bulk viscosity $\zeta(H)$ and the relaxation time $\tau(H)$ from a more complete microscopic theory of the phonon sector. One would like to start from an explicit effective field theory for the longitudinal modes of the brane or for additional scalar degrees of freedom, compute their dispersion relations and interaction rates, and evaluate the stress tensor correlator that enters the Kubo formula. This program would allow one to determine $\zeta(H)$ and $\tau(H)$ from first principles rather than from phenomenology and would test whether the simple scaling used here is robust or only a useful approximation. A more complete theory could clarify how the NG brane tension and the phonon fluid arise from a fundamental microscopic description and whether similar viscoelastic effects appear in other approaches to late time acceleration. The results presented here indicate that an elastic and viscous space is a viable candidate for the dark energy sector and they motivate a more systematic study of the microphysics and cosmological implications of such models.